# Visually Robust Adversarial Imitation Learning from Videos with Contrastive Learning

### v1

@@ Abstract
We propose C-LAIfO, a computationally efficient algorithm designed for imitation learning from videos, even in the presence of visual mismatch between agent and expert domains. We analyze the problem of imitation from expert videos with visual discrepancies, and introduce a solution for robust latent space estimation using contrastive learning and data augmentation. Provided a visually robust latent space, our algorithm performs imitation entirely within this space using off-policy adversarial imitation learning. We conduct a thorough ablation study to justify our design choices and test C-LAIfO on high-dimensional continuous robotic tasks. Additionally, we demonstrate how C-LAIfO can be combined with other reward signals to facilitate learning on a set of challenging hand manipulation tasks with sparse rewards. Our experiments show improved performance compared to baseline methods, highlighting the effectiveness and versatility of C-LAIfO. To ensure reproducibility, we provide open access to our code.
1 Introduction
@@
In this context, our paper introduces a novel end-to-end pipeline for imitation from expert videos with visual mismatch. We begin by analyzing the V-IfO problem with visual mismatch and propose a novel, simple, and computationally efficient algorithm called Contrastive Latent Adversarial Imitation from Observations (C-LAIfO). Notably, C-LAIfO builds upon the recent LAIfO algorithm [1] and achieves visually robust latent state estimation through data augmentation and contrastive learning techniques [3, 4]. We justify each design choice for our algorithm, including the types of data augmentation and contrastive loss, through a comprehensive ablation study. Furthermore, we evaluate the performance of C-LAIfO on the DeepMind Control Suite [5] and compare it against two V-IfO baselines—LAIfO [1] and PatchAIL [2]—as well as DisentanGAIL [6], which serves as a baseline for V-IfO with domain mismatch. Additionally, we show how the reward signal learned from expert data using C-LAIfO can be easily integrated with other signals to enhance efficiency and enable learning in scenarios with sparse rewards. In this context, we extend our algorithm’s evaluation to the Adroit platform for dynamic dexterous manipulation [7]. These additional experiments highlight the versatility of our approach, showcasing its efficacy in handling complex robotic tasks.
2 Related Work
Our work is closely related to the Adversarial Imitation Learning (AIL) framework [8], which has emerged as a promising approach to tackle the Imitation Learning (IL) problem. This framework builds upon a substantial body of research in inverse Reinforcement Learning (RL), where the primary objective is to identify a reward function that enables expert demonstrations (state-action pairs) to be optimal [9, 10, 11, 12, 13, 14]. The derived reward function is then used to train RL agents capable of achieving expert-level performance.
@@ 2 Related Work
AIL was initially formulated for the fully observable setting with access to expert actions [8, 15] and subsequently extended to observation-only scenarios in follow-up studies [16, 17, 18]. Furthermore, the stability and efficiency of this framework have been further investigated in [19, 20, 21, 22, 23]. In the context of partial observability, AIL has been explored for scenarios involving missing information [24], and in Visual IL where agents learn from video frames as state observations [25]. Relaxing the assumption on the availability of expert actions gives rise to the V-IfO problem, which is the primary focus of our work. State-of-the-art algorithms for the V-IfO setting include PatchAIL [2], which applies AIL directly on the pixel space utilizing a PatchGAN discriminator [26, 27], and LAIfO [1], where AIL operates on a latent representation of the agent state. Notably, these approaches are built on the assumption that both the expert and the learning agent act within the same decision process, which rarely holds true in real-world scenarios.
Our research focuses on the imitation problem from observations in the presence of domain mismatches between the expert and the learning agent. This framework is referred to in the literature under various names, including third-person IL [28], domain-adaptive IL [29], or cross-domain IL [30]. Previous solutions have approached this problem either by decomposing it into sequential stages [31] or using end-to-end methods. Among end-to-end approaches, some have addressed the fully observable setting [29, 32, 30, 33, 34], while others have tackled the visual (partially observable) setting [28, 35, 6, 36].
@@
In the visual setting, [28] and [6] propose methods to extract domain-independent features useful for inferring a domain-independent reward function. The work in [28] learns domain-independent features using an adversarial approach similar to [37], while DisentanGAIL in [6] achieves a similar result by adding a mutual information constraint to the binary cross-entropy loss used for AIL. Similar to our algorithm, these studies formulate fully end-to-end model-free algorithms that avoid costly generative steps during the imitation process. Our approach adopts similar reasoning to [28, 6]; however, we leverage contrastive learning for domain-independent feature extraction and build the entire AIL pipeline (both reward inference and RL step) on this learned feature space, rather than only the reward inference as done in [28, 6]. As shown in our experiments, this leads to significant improvements in performance.
In contrast, other works rely on expensive generative steps to address the mismatch problem. In [35], imitation is performed using expert observation-action pairs through a learned domain-agnostic recurrent state space model [38]. Our algorithm, on the other hand, is model-free and only requires expert observations. In [36], cycle-consistent adversarial networks [27] are used to generate expert videos in the agent’s domain, thus reducing the problem to the standard V-IfO without mismatches. Our approach does not require such a generative step, as it learns a domain-independent feature space directly.
3 Preliminaries
@@
Unless indicated otherwise, we use uppercase letters (e.g., $S_{t}$) for random variables, lowercase letters (e.g., $s_{t}$) for values of random variables, script letters (e.g., $\mathcal{S}$) for sets, and bold lowercase letters (e.g., $\bm{\theta}$) for vectors. Let $[t_{1}:t_{2}]$ be the set of integers $t$ such that $t_{1}\leq t\leq t_{2}$; we write $S_{t}$ such that $t_{1}\leq t\leq t_{2}$ as $S_{t_{1}:t_{2}}$. We denote with $\mathbb{E}[\cdot]$ expectation, with $\mathbb{P}(\cdot)$ probability, and with $\mathbb{D}_{f}(\cdot,\cdot)$ an $f$-divergence between two distributions of which the Jensen-Shannon divergence, $\mathbb{D}_{\text{JS}}(\cdot||\cdot)$, is a special case.
@@
Maximizing (1) with respect to $\bm{\chi}$ is effectively an inverse RL step where a reward function, $r_{\bm{\chi}}(s,a)=-\log(1-D_{\bm{\chi}}(s,a))$, is inferred by leveraging $\tau_{E}$ and $\tau_{\bm{\theta}}$. On the other hand, minimizing (1) with respect to $\bm{\theta}$ can be interpreted as an RL step, where the agent aims to minimize its expected cost. It has been demonstrated that optimizing the min-max objective in (1) is equivalent to minimizing $\mathbb{D}_{\text{JS}}(\rho_{\pi_{\bm{\theta}}}(s,a)||\rho_{\pi_{E}}(s,a))$, so we are recovering the expert state-action visitation distribution [39].
@@
4 Contrastive Latent Adversarial Imitation from Observations
@@
In the following, we present the main components of our algorithm C-LAIfO, which performs imitation directly in a domain-invariant feature space $\mathcal{Z}$. In order to do so, we learn a domain-invariant encoder $\phi_{\bm{\delta}}$ that can successfully map $x^{T}_{\leq t}$ and $x^{S}_{\leq t}$ to $z_{t}$ through two main steps. First, we train the encoder $\phi_{\bm{\delta}}$ alongside the critic networks $Q_{\bm{\psi}_{k}}$ where $k=\{1,2\}$. This step is essential for solving the imitation problem and embedding goal-completion information within the latent space $\mathcal{Z}$. We further train $\phi_{\bm{\delta}}$ to optimize an auxiliary contrastive loss and perform randomized augmentation of the observations, taking into account the type of visual mismatch between the source and the target domains. This step is crucial to efficiently discard visually-distracting information from $\mathcal{Z}$.
Adversarial imitation in latent space
@@
$\max_{\bm{\chi}}\ \ \mathbb{E}_{(z_{\bm{\delta}},z_{\bm{\delta}}^{\prime})\sim\mathcal{B}_{E}}[\log(D_{\bm{\chi}}(z_{\bm{\delta}},z_{\bm{\delta}}^{\prime}))]+\mathbb{E}_{(z_{\bm{\delta}},z_{\bm{\delta}}^{\prime})\sim\mathcal{B}}[\log(1-D_{\bm{\chi}}(z_{\bm{\delta}},z_{\bm{\delta}}^{\prime}))].$ (2)
@@
Critic and encoder training step
@@
$\min_{\bm{\psi}_{k},\bm{\delta}}$ $\ \ \ \mathbb{E}_{(\tilde{z}_{\bm{\delta}},a_{t},\tilde{z}_{\bm{\delta}}^{\prime})\sim\mathcal{B}}[\big{(}Q_{\bm{\psi}_{k}}(\tilde{z}_{\bm{\delta}},a_{t})-\text{sg}(y)\big{)}^{2}]+\mathbb{E}_{\tilde{z}_{\bm{\delta}}\sim\mathcal{B}}[\mathcal{L}(\tilde{z}_{\bm{\delta}})]$ (3)
@@
The steps in (3)–(4) follow the deep $Q$-network optimization pipeline [41, 42], where we add a contrastive auxiliary loss $\mathbb{E}_{\tilde{z}_{\bm{\delta}}\sim\mathcal{B}}[\mathcal{L}(\tilde{z}_{\bm{\delta}})]$ in (3) defined on the encoder $\phi_{\bm{\delta}}$. In (4), the reward function $r_{\bm{\chi}}(z_{\bm{\delta}},z_{\bm{\delta}}^{\prime})$ is computed through the AIL step in (2). Note that we do not perform data augmentation when computing $r_{\bm{\chi}}$ since, during reward inference, we are deploying and not training the encoder $\phi_{\bm{\delta}}$. In practice, adding data augmentation to AIL in (2) decreases the performance.
In (3), the encoder $\phi_{\bm{\delta}}$ is trained together with the critic networks $Q_{\bm{\psi}_{k}}$ ($k=\{1,2\}$) in order to regress $y$ in (4), where $\text{sg}(\cdot)$ stands for stop gradient of the encoder parameters $\bm{\delta}$. The value of $y$ in (4) is computed by summing $r_{\bm{\chi}}$ with the discounted target critic network at time $t+1$. In (4), $a^{\prime}=\pi_{\bm{\theta}}(\tilde{z}_{\bm{\delta}}^{\prime})+\epsilon$ where $\epsilon\sim\text{clip}(\mathcal{N}(0,\sigma^{2}),-c,c)$ is a clipped exploration noise with $c$ the clipping parameter and $\mathcal{N}(0,\sigma^{2})$ a univariate normal distribution with zero mean and $\sigma$ standard deviation. $\bar{\bm{\psi}}_{1}$ and $\bar{\bm{\psi}}_{2}$ are the slow moving weights for the target critic networks. $\mathcal{B}$ is a replay buffer initialized to store interactions $(x^{T}_{t},a_{t},x^{T}_{t+1})$ of the agent with the target environment. Note that the latent state-transitions $(\tilde{z}_{\bm{\delta}},\tilde{z}_{\bm{\delta}}^{\prime})$ are inferred from sequences of observations using $\phi_{\bm{\delta}}$ and, as above, we write $(\tilde{z}_{\bm{\delta}},\tilde{z}_{\bm{\delta}}^{\prime})\sim\mathcal{B}$ to streamline the notation (cf. Appendix B for the complete pseudo-code).
@@
Contrastive loss
In the following, we introduce the data augmentation techniques and the auxiliary loss $\mathcal{L}$ in (3) for C-LAIfO. We opt for a contrastive method as this leads to good empirical results and good computational efficiency. Contrastive learning constructs low-dimensional representations of high-dimensional data by maximizing agreement between augmented views of the same data example via a contrastive loss in the latent space $\mathcal{Z}$. In our specific case, we define equivalent data as sequences of observations with the same goal-completion information. The data augmentation is randomized by considering a set of pre-determined functions. We will show in our experiments (Section 5) that both the choice of contrastive loss and the set of augmentation functions play an important role in filtering out visually-distracting information.
@@
$\mathcal{L}(z_{\bm{\delta}})=-\log\frac{\exp(\text{sim}(\tilde{z}_{\bm{\delta}}(i),\tilde{z}_{\bm{\delta}}(j))/\eta)}{\sum_{k=1}^{2N}\mathbbm{1}_{[k\neq i]}\exp(\text{sim}(\tilde{z}_{\bm{\delta}}(i),\tilde{z}_{\bm{\delta}}(k))/\eta)},$ (5)
where $\mathbbm{1}_{[k\neq i]}\in\{0,1\}$ is an indicator function equal to 1 if $k\neq i$, $\eta$ denotes a temperature parameter, and $\text{sim}(\bm{u},\bm{v})=\bm{u}^{\intercal}\bm{v}/||\bm{u}||||\bm{v}||$ denotes the cosine similarity. We sample a batch of $N$ sequences of observations from the buffer $\mathcal{B}$ and define pairs of augmented sequences derived from the batch, resulting in $2N$ data points. The negative data points are not sampled explicitly. Instead, given a positive pair, we treat the other $2(N-1)$ augmented data points within the batch as negatives. Note that the loss in (5) is called the normalized temperature-scaled cross entropy loss or the Information Noise-Contrastive Estimation (InfoNCE) loss [43] and represents an upper bound of the negative mutual information between positive pairs. Therefore, by minimizing (5) as in (3) we are maximizing the mutual information between positive pairs in the latent space $\mathcal{Z}$.
5 Experiments
In this section, we present our experimental results. We begin with an ablation study to justify the key design choices of our algorithm. Next, we demonstrate how C-LAIfO effectively handles various types of visual mismatches in the V-IfO setting. Finally, we showcase how C-LAIfO facilitates learning in challenging robotic manipulation tasks with sparse rewards and realistic visual inputs.
Ablation study
Given the environments in Fig. 2, we perform the following ablations:

### v2

@@ Abstract
We propose C-LAIfO, a computationally efficient algorithm designed for imitation learning from videos in the presence of visual mismatch between agent and expert domains. We analyze the problem of imitation from expert videos with visual discrepancies, and introduce a solution for robust latent space estimation using contrastive learning and data augmentation. Provided a visually robust latent space, our algorithm performs imitation entirely within this space using off-policy adversarial imitation learning. We conduct a thorough ablation study to justify our design and test C-LAIfO on high-dimensional continuous robotic tasks. Additionally, we demonstrate how C-LAIfO can be combined with other reward signals to facilitate learning on a set of challenging hand manipulation tasks with sparse rewards. Our experiments show improved performance compared to baseline methods, highlighting the effectiveness of C-LAIfO. To ensure reproducibility, we open source our code.
I INTRODUCTION
@@
We introduce a novel end-to-end pipeline for imitation from expert videos with visual mismatch. We begin by analyzing the V-IfO problem with visual mismatch and propose a novel, simple, and computationally efficient algorithm called Contrastive Latent Adversarial Imitation from Observations (C-LAIfO). Notably, C-LAIfO builds upon the recent LAIfO algorithm [1] and achieves visually robust latent state estimation through data augmentation and contrastive learning techniques [3, 4]. We justify each design choice for our algorithm, including the types of data augmentation and contrastive loss, through a comprehensive ablation study. Furthermore, we compare C-LAIfO against two V-IfO baselines—LAIfO [1] and PatchAIL [2]—as well as DisentanGAIL [5], which serves as a baseline for V-IfO with domain mismatch. Additionally, we show how the reward signal learned from expert data using C-LAIfO can be easily integrated with other signals to enhance efficiency and enable learning in robotic tasks with sparse reward functions. Therefore, we further evaluate our algorithm on the Adroit platform for dynamic dexterous manipulation [6]. These additional experiments highlight the versatility of our approach, showcasing its efficacy in handling complex robotic tasks.
II Related Work
Imitation from observation
@@ II Related Work
Imitation Learning (IL) is a powerful approach that allows agents to mimic expert behavior by using demonstrations of a task typically in the form of state-action pairs. Our work builds on Adversarial Imitation Learning (AIL) [7, 8] which frames IL as a two-player game between a discriminator and an agent’s policy. Here, the discriminator distinguishes whether a state-action pair is generated by the agent or the expert policy. In practice, AIL is formulated as a joint process of Reinforcement Learning (RL) and inverse RL [9, 10, 11, 12]. First a reward function is inferred from expert demonstrations and then it is used in the RL step to train agents. In scenarios with partial observability, AIL has been applied to cases with missing information [13] and to Visual IL, where agents learn from video frames as state observations [14]. Compared to standard IL, Imitation from observation [15, 16, 17] assumes that action information is not observable in the demonstrations data. This setting is more practical compared to IL but also more difficult to tackle. The combination of learning from videos in the absence of expert actions gives rise to the V-IfO problem, which is the primary focus of our work. End-to-end state-of-the-art algorithms for the V-IfO setting include PatchAIL [2], which applies AIL directly on the pixel space utilizing a PatchGAN discriminator [18, 19], and LAIfO [1], where AIL operates on a latent representation of the agent state. Notably, these approaches are built on the assumption that both the expert and the learning agent act within the same decision process, which rarely holds true in real-world scenarios.
Imitation from videos with environment mismatch
Our research targets imitation from visual observations in the presence of mismatches between the expert and the learner environments, a problem referred to as third-person IL [20], domain-adaptive IL [21], or cross-domain IL [22]. Solutions in the literature either decompose this problem into sequential stages or formulate end-to-end methods. Sequential approaches include: [23], where a reward function is learned by leveraging video prediction with context translation; [24], where reward functions are obtained using time-contrastive networks trained offline; [25], where cycle-consistent adversarial networks [19] are trained offline to generate instruction images in the agent domain from videos; [26], which uses inverse models and adversarial domain adaptation [27] to train navigation policies from videos, and [28], where 3D trajectory reconstruction from videos is used to obtain physically plausible trajectories. Our work differs from this literature as we formulate a fully end-to-end approach.
End-to-end algorithms for imitation from videos with mismatch
@@
Previous end-to-end solutions were presented in [20, 29, 5, 30]. The studies in [20] and [5] extract domain-independent features to infer domain-independent reward functions. More specifically, in [20], the authors propose to learn domain-independent features using an adversarial approach similar to [31], while DisentanGAIL in [5] achieves a similar result by adding a mutual information constraint to the binary cross-entropy loss used for AIL. Similar to our algorithm, these studies formulate fully end-to-end model-free algorithms that avoid costly generative steps during the imitation process. Our approach adopts similar reasoning to [20, 5]; however, we leverage contrastive learning for domain-independent feature extraction and build the entire AIL pipeline (both reward inference and RL step) on this learned feature space, rather than only the reward inference as done in [20, 5]. As shown in our experiments, this leads to significant improvements in performance. Other works rely on expensive generative steps to address the mismatch problem. In [29], imitation is performed using expert observation-action pairs through a learned domain-agnostic recurrent state space model [32]. Our algorithm, on the other hand, is model-free and only requires expert observations. In [30], cycle-consistent adversarial networks [19] are trained online to generate expert videos in the agent’s domain, thus reducing the problem to the standard V-IfO without mismatches. Our approach does not require such a generative step, as it learns a domain-independent feature space directly.
III Preliminaries
@@
We use uppercase letters (e.g., $S_{t}$) for random variables, lowercase letters (e.g., $s_{t}$) for values of random variables, script letters (e.g., $\mathcal{S}$) for sets, and bold lowercase letters (e.g., $\bm{\theta}$) for vectors. Let $[t_{1}:t_{2}]$ be the set of integers $t$ such that $t_{1}\leq t\leq t_{2}$; we write $S_{t}$ such that $t_{1}\leq t\leq t_{2}$ as $S_{t_{1}:t_{2}}$. We denote with $\mathbb{E}[\cdot]$ expectation, with $\mathbb{P}(\cdot)$ probability, and with $\mathbb{D}_{f}(\cdot,\cdot)$ an $f$-divergence between two distributions of which the Jensen-Shannon divergence, $\mathbb{D}_{\text{JS}}(\cdot||\cdot)$, is a special case.
@@
Maximizing (1) with respect to $\bm{\chi}$ is effectively an inverse RL step where a reward function, $r_{\bm{\chi}}(s,a)=-\log(1-D_{\bm{\chi}}(s,a))$, is inferred by leveraging $\tau_{E}$ and $\tau_{\bm{\theta}}$. Minimizing (1) with respect to $\bm{\theta}$ is an RL step, where the agent aims to minimize its expected cost. Optimizing (1) is equivalent to minimizing $\mathbb{D}_{\text{JS}}(\rho_{\pi_{\bm{\theta}}}(s,a)||\rho_{\pi_{E}}(s,a))$, so we are recovering the expert state-action visitation distribution [33].
@@
IV Contrastive Latent Adversarial Imitation from Observations
Figure 2: Summary of C-LAIfO. In the diagram, black lines indicate shared weights among networks, blue arrows indicate forward pass through the networks, and red arrows indicate backward pass. The losses $\mathcal{L}_{D}$, $\mathcal{L}_{Q}$ and $\mathcal{L}(z_{\bm{\delta}})$ are respectively in (2), (3), and (5). $\mathcal{L}_{\pi}$ indicates the deterministic actor-critic loss [34].
@@ IV Contrastive Latent Adversarial Imitation from Observations
In the following, we present the main components of our algorithm C-LAIfO, which performs imitation directly in a domain-invariant feature space $\mathcal{Z}$ (Sec. IV-A). In order to do so, we learn a domain-invariant encoder $\phi_{\bm{\delta}}$ that can successfully map $x^{T}_{\leq t}$ and $x^{S}_{\leq t}$ to $z_{t}$ through two main steps. First, we train the encoder $\phi_{\bm{\delta}}$ alongside the critic networks $Q_{\bm{\psi}_{k}}$ where $k=\{1,2\}$ (Sec. IV-B). This step is essential for solving the imitation problem and embedding goal-completion information within the latent space $\mathcal{Z}$. We further train $\phi_{\bm{\delta}}$ to optimize an auxiliary contrastive loss and perform randomized augmentation of the observations, taking into account the type of visual mismatch between the source and the target domains (Sec. IV-C). This step is crucial to efficiently discard visually-distracting information from $\mathcal{Z}$. A schematic diagram summarizing the whole C-LAIfO pipeline is provided in Fig. 2.
IV-A Adversarial imitation in latent space
@@
$\begin{split}\max_{\bm{\chi}}\ &\mathbb{E}_{(z_{\bm{\delta}},z_{\bm{\delta}}^{\prime})\sim\mathcal{B}_{E}}[\log(D_{\bm{\chi}}(z_{\bm{\delta}},z_{\bm{\delta}}^{\prime}))]\\ &+\mathbb{E}_{(z_{\bm{\delta}},z_{\bm{\delta}}^{\prime})\sim\mathcal{B}}[\log(1-D_{\bm{\chi}}(z_{\bm{\delta}},z_{\bm{\delta}}^{\prime}))].\end{split}$ (2)
@@
IV-B Critic and encoder training step
@@
$\begin{split}\min_{\bm{\psi}_{k},\bm{\delta}}&\ \mathbb{E}_{(\tilde{z}_{\bm{\delta}},a_{t},\tilde{z}_{\bm{\delta}}^{\prime})\sim\mathcal{B}}[\big{(}Q_{\bm{\psi}_{k}}(\tilde{z}_{\bm{\delta}},a_{t})-\text{sg}(y)\big{)}^{2}]\\ &\ \ \ +\mathbb{E}_{\tilde{z}_{\bm{\delta}}\sim\mathcal{B}}[\mathcal{L}(\tilde{z}_{\bm{\delta}})]\end{split}$ (3)
@@
The steps in (3)–(4) follow the deep $Q$-network optimization pipeline [36, 37], where we add a contrastive auxiliary loss $\mathbb{E}_{\tilde{z}_{\bm{\delta}}\sim\mathcal{B}}[\mathcal{L}(\tilde{z}_{\bm{\delta}})]$ in (3) defined on the encoder $\phi_{\bm{\delta}}$. In (4), the reward function $r_{\bm{\chi}}(z_{\bm{\delta}},z_{\bm{\delta}}^{\prime})$ is computed through the AIL step in (2). Note that we do not perform data augmentation when computing $r_{\bm{\chi}}$ since, during reward inference, we are deploying and not training the encoder $\phi_{\bm{\delta}}$. In practice, adding data augmentation to AIL in (2) decreases the performance. Refer to the Arxiv version for empirical evidence on this claim. In (3), the encoder $\phi_{\bm{\delta}}$ is trained together with the critic networks $Q_{\bm{\psi}_{k}}$ ($k=\{1,2\}$) in order to regress $y$ in (4), where $\text{sg}(\cdot)$ stands for stop gradient of the encoder parameters $\bm{\delta}$. The value of $y$ in (4) is computed by summing $r_{\bm{\chi}}$ with the discounted target critic network at time $t+1$. In (4), $a^{\prime}=\pi_{\bm{\theta}}(\tilde{z}_{\bm{\delta}}^{\prime})+\epsilon$ where $\epsilon\sim\text{clip}(\mathcal{N}(0,\sigma^{2}),-c,c)$ is a clipped exploration noise with $c$ the clipping parameter and $\mathcal{N}(0,\sigma^{2})$ a univariate normal distribution with zero mean and $\sigma$ standard deviation. $\bar{\bm{\psi}}_{1}$ and $\bar{\bm{\psi}}_{2}$ are the slow moving weights for the target critic networks. $\mathcal{B}$ is a replay buffer initialized to store interactions $(x^{T}_{t},a_{t},x^{T}_{t+1})$ of the agent with the target environment. Note that the latent state-transitions $(\tilde{z}_{\bm{\delta}},\tilde{z}_{\bm{\delta}}^{\prime})$ are inferred from sequences of observations using $\phi_{\bm{\delta}}$ and, as above, we write $(\tilde{z}_{\bm{\delta}},\tilde{z}_{\bm{\delta}}^{\prime})\sim\mathcal{B}$ to streamline the notation. Refer to the Arxiv version for the complete pseudo-code.
@@
IV-C Contrastive loss
In the following, we introduce the data augmentation techniques and the auxiliary loss $\mathcal{L}$ in (3) for C-LAIfO. We opt for a contrastive method as this leads to good empirical results and good computational efficiency. Contrastive learning constructs low-dimensional representations of high-dimensional data by maximizing agreement between augmented views of the same data example via a contrastive loss in the latent space $\mathcal{Z}$. In our specific case, we define equivalent data as sequences of observations with the same goal-completion information. The data augmentation is randomized by considering a set of pre-determined functions. We will show in our experiments (Section V) that both the choice of contrastive loss and the set of augmentation functions play an important role in filtering out visually-distracting information.
@@ IV-C Contrastive loss
$\mathcal{L}(z_{\bm{\delta}})=-\log\frac{\exp(\text{sim}(\tilde{z}_{\bm{\delta}}(i),\tilde{z}_{\bm{\delta}}(j))/\eta)}{\sum_{k=1}^{2N}\mathds{1}_{[k\neq i]}\exp(\text{sim}(\tilde{z}_{\bm{\delta}}(i),\tilde{z}_{\bm{\delta}}(k))/\eta)},$ (5)
where $\mathds{1}_{[k\neq i]}\in\{0,1\}$ is an indicator function equal to 1 if $k\neq i$, $\eta$ denotes a temperature parameter, and $\text{sim}(\bm{u},\bm{v})=\bm{u}^{\intercal}\bm{v}/||\bm{u}||||\bm{v}||$ denotes the cosine similarity. We sample a batch of $N$ sequences of observations from the buffer $\mathcal{B}$ and define pairs of augmented sequences derived from the batch, resulting in $2N$ data points. The negative data points are not sampled explicitly. Instead, given a positive pair, we treat the other $2(N-1)$ augmented data points within the batch as negatives. Note that the loss in (5) is called the normalized temperature-scaled cross entropy loss or the Information Noise-Contrastive Estimation (InfoNCE) loss [38] and represents an upper bound of the negative mutual information between positive pairs. Therefore, by minimizing (5) as in (3) we are maximizing the mutual information between positive pairs in the latent space $\mathcal{Z}$.
V Experiments
In this section, we begin with an ablation study to justify the design choices of our algorithm (Sec. V-A). Next, we demonstrate how C-LAIfO effectively handles various types of visual mismatches in the V-IfO setting (Sec. V-B). Finally, we showcase how C-LAIfO facilitates learning in challenging robotic manipulation tasks with sparse rewards and realistic visual inputs (Sec. V-C). In all the experiments, we use DDPG [39] to train experts in a fully observable setting and collect $100$ episodes of expert data. The learned policies are evaluated based on average return over $10$ episodes. We report the mean and standard deviation of the final return over $6$ seeds and highlight the best performance.
V-A Ablation study
TABLE I: Summary of the ablation experiments. All algorithms are trained for $10^{6}$ steps. The experiments C-LAIfO w/o $Q$ backprop, C-LAIfO full aug, and C-LAIfO w/o aug are only conducted in the easier setting due to their low performance.
Source Env: [Uncaptioned image], Performance = $950$
Target Env [Uncaptioned image] [Uncaptioned image]
C-LAIfO $808\pm 269$ $768\pm 231$
BYOL-LAIfO $707\pm 337$ $142\pm 124$
C-LAIfO w/o $Q$ backprop $48.7\pm 8.7$
C-LAIfO full aug $96.8\pm 53.4$
C-LAIfO w/o aug $113\pm 25.3$
In this section, we perform the following ablations: